The overpic environment with relative positioning^{1}

Insert before:

\usepackage[percent]{overpic}
(default) or
\usepackage[permil]{overpic}

The original graphic from GhostScript:

\includegraphics[scale=.25]{golfer.ps}

0 10 20 30 40 50 60 70 80 90 100 0 10 20 30 40 50 60 70 80 90 100 0 10 20 30 40 50 60 70 80 90 100 0 10 20 30 40 50 60 70 80 90 100

A grid for orientation:

\begin{overpic}[scale=.25,grid,tics=10]%
      {golfer.ps}
\end{overpic}

(default for tics is 10 for ‘percent’ and 100 for

‘permil’)

L^{A}TEX

Positioning the L^{A}TEX output with \put commands:

\begin{overpic}[scale=.25]{golfer.ps}
  \put(5,45){\huge \LaTeX}
  \put(55,10){\includegraphics[scale=.07]%
      {golfer.ps}}
\end{overpic}

(\unitlength is internally set to 1/100 or 1/1000 of the length of the picture’s longer side)

1Version 0.5, Rolf Niepraschk, 11/97, file: overpic.sty